Reece:

The sum of two numbers is 45. Their difference is 11. What is the smaller number?

Reece:

help plzzz

DumbFounded:

ok so

DumbFounded:

let the numbers be a and b so a + b = 45 and a - b = 11

DumbFounded:

what do we get when we add the numbers?

Reece:

56?

DumbFounded:

yes good

Reece:

is that the answer

DumbFounded:

no

DumbFounded:

whats the answer for a - b = 11?

Reece:

Traveling with the current, a certain boat went 15 km/h. Against the same current, it only went 1 km/h. Find the speed of the boat in still water and the speed of the current. The speed of the boat is? The speed of the current is?

Reece:

what about this one

DumbFounded:

lets stick with the first problem

DumbFounded:

im not allowed to give u the answers i have to guide u through them

DumbFounded:

whats a - b = 11?

Reece:

but i jus

DumbFounded:

what?

DumbFounded:

wait let me show u some of the equations

Reece:

idk whatsa-b=11

DumbFounded:

Let the numbers be a and b Given that a+b=45 ....(1) a−b=11 ....(2) Adding these two equations, we get 2a=56 ⇒a=28

DumbFounded:

now you have to substitute

DumbFounded:

28+b=45

DumbFounded:

ty @jhonyy9

Reece:

so we gotta do 28+b=45

DumbFounded:

yes

DumbFounded:

b=45-28 right?

Reece:

yes

DumbFounded:

since we substituted

DumbFounded:

and b also equals 17

Reece:

right

DumbFounded:

so im going to show you all the steps now

Reece:

ok

DumbFounded:

Let the numbers be a and b Given that a+b=45 ....(1) a−b=11 ....(2) Adding these two equations, we get 2a=56 ⇒a=28 Substituting value of a in equation (1), we get 28+b=45 ⇒b=45−28 ⇒b=17 We get a=28 and b=17

Reece:

is that for the firstproblem i showed you??

DumbFounded:

do u understand now?

DumbFounded:

yes

Reece:

and the answer woud be17

DumbFounded:

now u want me to solve the second problem?

Reece:

plz

DumbFounded:

ok so

DumbFounded:

let v = the rate of the boat in still water c = the rate of the current d = the distance traveled in one direction

DumbFounded:

the canoe's speed downstream: v + c = 11 km/h the canoe's speed upstream: v - c = 3 km/h

DumbFounded:

right?

Reece:

yes

DumbFounded:

wait wait wait im sorry i have the wrong problem

Reece:

ohh

DumbFounded:

ok now im stuck lol

DumbFounded:

@jhonyy9 can you help him with his second problem?

DumbFounded:

jhonny will be with you in a minute

Reece:

waiti have a question for number one

DumbFounded:

ok shoot

Reece:

what is the smaller number for number 1

Reece:

The sum of two numbers is 45. Their difference is 11. What is the smaller number?

Reece:

what about this problem When I looked outside, I saw birds and cats. I counted 19 heads and 42 legs. ( I counted because I thought: This would be a good math problem!) How many birds and how many cats did I see in the yard

Reece:

@Vocaloid r u good at math??

ActRight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Reece and the answer woud be17 \(\color{#0cbb34}{\text{End of Quote}}\) @Angle

Reece:

????